OpenStudy (anonymous):

Hunus: A rental company charges $26.50 to rent a wheelchair for the first week and $21.50 for every week after the first. If Sarah's rental fee was $69.50, which equation could be used to find out how many weeks (w) she used the wheel chair?

OpenStudy (anonymous):

You don't have to simplify it :) and again, thank you so much. I am really grateful!

OpenStudy (hunus):

\[26.50+21.50x=69.50\] \[21.50x=43\] \[x=2\]

OpenStudy (anonymous):

There's D. 21.5(w-1)+26.5=69.5 B. 26.5w-21.5=69.5 and then there's E. 21.5(w-1)=69.5 but no 26.5+21.5x=69.5

OpenStudy (hunus):

D. 21.5(w-1)+26.5=69.5

OpenStudy (hunus):

Sorry lol. Forgot to count 26.50 as the first week